# Quantum Imaging and Metrology with Undetected squeezed Photons: Noise Canceling and Noise Based Imaging

### v1

@@ Abstract
In this work a quantum imaging setup based on undetected squeezed photons is employed for metrological applications such as sensitive phase measurement and quantum imaging. In spite of the traditional quantum imaging with undetected photons, introduced by A. Zeilinger et. al, the proposed setup is equipped by a homodyne detection and also the brightness of the quantum light is enhanced by an optical parametric oscillator (OPO). Introducing OPO may be diminish the validity of the low gain approximation, so a theoretical approach beyond this approximation is introduced. Due to the resource of squeezing, the results reveal the higher amount of signal to noise ratio, as a measure of image quality and phase-measurement accuracy. Accordingly, an imaging protocol is introduced to suppress the background noises, effectively. Interestingly, This protocol provides a way to extract the image information which is encoded in the quantum fluctuation (noise). Therefore, non-disruptive imaging is achievable, which is noteworthy subject in the field of bio-imaging of sensitive and low damage threshold living cells.
I Introduction
Interferometery provides a powerful technique to measure physical quantities with high precision and sensitivity[1]. In the conventional interferometers the sensitivity is bounded by Shot-Noise-Limit (SNL)[2, 3, 4]. The ultimate goal in the interferometery is to overcome SNL. Non-classical properties of lights may provide an effective resources to gain this purpose[5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. In recent years, along with development of non-linear optical process, the optical non-linear interferometery attracted extensive attentions, to realize interferometer with strong sensitivity and accuracy [16, 17, 18, 19]. SU(1,1) interferometers are almost the most important non-linear interferometer due to their sensitivity enhancement[20, 16, 21, 22, 23, 24, 25, 26]. These interferometers are used in direct detection of gravitational waves[25].
In the SU(1,1) interferometer, the beam splitters of conventional interferometer are replaced by a two-mode squeezer apparatuses such as Optical Parametric Amplifiers(OPAs)[27]. In the devices, which are using non-degenerate SU(1,1) interferometer, the interaction and detection often applied in different wavelength[28, 29, 30]. In the other word, these type of device can provide two mode squeezed light with different wavelength called signal and idler beams. Operating in two non-degenerate wavelength makes SU(1,1) interferometer to become powerful devices in quantum metrology such as spectroscopy[31, 32], imaging[33, 34] and optical coherent tomography(OCT)[2, 35]. This properties have significant application in image construction of biological molecules and living cells[36, 37] because the interaction could be occur in mid-infrared and detection is applied in the optical spectrum[38].
Quantum Imaging with Undetected Photon (QUIP) is an imaging protocol in which SU(1,1) interferometer is used in image formation such that the idler beam interacts with an object while detection of the object achieved by signal beam which never interacts with the object[30]. This type of imaging is developed by Anton Zeilinger et al.[39]. In their setup, a pair signal and idler beam are generated by Spontaneous Emission Down Conversion (SPDC)process in a non-linear crystal. The idler beam guided to the second non-linear crystal after interacting with the object. In the second non-linear crystal a new signal and idler pair are created. The signal generated by the first and second non-linear crystals interfer to form the image. The construction and deconstruction patterns appears due to the path indistinguishablity which is related to the quantum nature of the light[40]. The quantum theory of QUIP is developed by Mayukh Lahiri et. al. in Ref.[41]. In this research the low gain regime is used to model wave mixing in the non-linear crystals. The resolution of QUIP is studied in near-field and far-field imaging using Gaussian beam profiles[42]. In the near-field imaging the minimum resolvable distance between two point sources is determined by Point Spread Function (SPF) width which is proportional to the square root of crystal length[43], while in the far-field QUIP resolution is independent to the length of the non-linear crystals[43]. Moreover QUIP protocol is implemented in the microscopy[44] and holography devices[45].
However, since in the theoretical treatment of QUIP the low gain approximation is used, the results cannot be applied to the setup for those the low gain approximation is invalid. In this paper we present the theoretical treatment of QUIP beyond low gain approximation. To this end, it assumed that the non-linear crystals are derived by classical pump beam. Since in the present theoretical description the low gain approximation is ignored, it is applicable to the QUIP with desired gain regime. It is also, we introduce an imaging protocol in which one can suppress the background noises. The novel scheme is the image construction from a low damage threshold object which is known as Quadrature-Noise Shadow Imaging (QSI). This scheme is based on quadrature quantum noise measurement[46]. QSI allows to image with low photons number. Moreover, in the QSI the detrimental dark count noise of the sensor is eliminated in the real setups [46, 47, 48]. Interestingly, we found that the image information is also coded into the noise by which the image may reconstructed using QSI.
@@ I Introduction
This paper is organized as follow: In the next section the setup under consideration is described. Section III is devoted to the theoretical description of the model. In first subsection of Sec.IV the object low transmitivity is characterized while in second subsection the high transmitivity is addressed. Accordingly two imaging protocols are introduced in Sec. V. Finally the conclusion and remarks of this work are provided in Sec. VI.
II Proposed Setup
@@
In this section we demonstrate the proposed setup. The schematic diagram of the setup is depicted in figure 1. The image reconstruction is shown in figure 1 through four stages. In the first stage the two mode squeezed state beams are generated via a non-degenerated Optical Parametric Oscillator (NONPO), constructed by a non-linear crystal ($\chi_{1}^{(2)}\neq 0$) implemented in the center of a confocal cavity. Then each squeezed mode (output of coupler $M_{2}$ ) is separated to signal and idler mode by a dichoric mirror, $DM$. The signal (green) beam arrives to the second non-linear crystal ($\chi_{2}^{(2)}\neq 0$) and the idler (red) beam is reflected through $DM$ toward the object. The idler interacts with an object modeled by a lossless beam splitter with transmission $Te^{i\phi_{T}}$ (reflection $Re^{i\phi_{R}}$). In order to provide path identity, a portion of the idler, which is transmitted through the object, is combined with the idler beam generated in the second non-linear crystal. The information about the object is transferred from idlers to signals beam due to the three wave mixing phenomenon which occurs in the 2nd crystal, $\chi^{(2)}_{2}$. Finally, since the object information are encoded in the signal quadrature, a balance homodyne detection is applied to the outcome signal of the second non-linear crystal to extract the image information. It is worth noting that undetected light refers to the idler beam as no measurement applied to it.
@@
In the following section we provide a theoretical description of how the image information extracted.
III Theoretical Description of The Model
In this section we present theoretical description about image formation. As shown in figure 1, the output electrical signal of the balanced homodyne detection can be calculated as:
@@ III Theoretical Description of The Model
where we assumed that the local oscillator in the homodyne detection is prepared in coherent state $\left|\beta e^{i\phi_{\beta}}\right>$. The annihilation (creation) operator $a_{S2}$ ($a_{S2}^{\dagger}$) is correspond to the output signal beam from second non-linear crystal. As mentioned in the previous section, the object information is encoded into beam, $S2$. So, the aim is to relate $a_{S2}$ to the initial condition and object properties. To this end we assume that NOPO is derived by classical field with fixed phase, $\phi_{P1}$. The classical pump field implies the wave mixing process in the first non-linear crystal, $\chi_{1}^{(2)}$. Since the squeezing occurs via parametric down conversion process, the output of NOPO becomes,
@@
$a_{IT}=Te^{i\phi_{T}}a_{I1}+Re^{i\phi_{R}}a_{IL}.$ (3)
In this equation, the input field to empty port of the beam splitter (object) is denoted by $a_{IL}$. Using Eq. (2b) in (3), gives
@@
Since two beams ($a_{S1}$ and $a_{IT}$) are mixed in the second non-linear crystal and gives rise to built up beam $S_{2}$, we have,
@@
where the non-linear (squeezing) process in the second crystal is characterized by $G_{2}$, $g_{2}$ and $\phi_{P2}$ is the phase of classical field which pumps the second wave mixing process. Substituting Eqs. (2a) and (4) into Eq. (5) leads to,
$a_{S2}=Ga_{S0}+ga_{I0}^{\dagger}+ra_{IL}^{\dagger}.$ (6)
@@
where $G=G_{1}G_{2}+g_{1}g_{2}T\exp\left(i\phi_{1}\right)$, $g=g_{1}G_{2}\exp(i\phi_{P1})+G_{1}g_{2}T\exp\left(i\phi_{2}\right)$ and $r=g_{2}R\exp(i\phi_{3})$. In the expression for $G$, $g$ and $r$ the definitions $\phi_{1}=\phi_{P2}-\phi_{P1}-\phi_{T}$, $\phi_{2}=\phi_{P2}-\phi_{T}$ and $\phi_{3}=\phi_{P2}-\phi_{R}$ are used. It is worth noting, since $|G|^{2}-|g|^{2}-|r|^{2}=1$ is satisfied, the commutation relation $\left[a_{S2},a_{S2}^{\dagger}\right]=1$ is also preserved.
IV Description Characteristic of the System
[…]
$\langle S\left(\delta,\Delta\right)\rangle=2\alpha\beta|G|\cos\left(\phi_{G}(\Delta)+\delta\right),$ (7)
where, $\delta=\phi_{\alpha}+\phi_{\beta}$, $\Delta=\phi_{P2}-\phi_{P1}$ and $\phi_{G}$ is defined as,
$\phi_{G}(\Delta)=\tan^{-1}\left(\frac{x\sin(\Delta-\phi_{T})}{1+x\cos(\Delta-\phi_{T})}\right),$ (8)
where $x=Tg_{1}g_{2}(G_{1}G_{2})^{-1}$. Since $g_{i}$ and $G_{i}$ ($i=1,2$) are correspond to the parametric process gains and $T\leq 1$, the condition $0\leq x\leq 1$ is satisfied. In Eq. (7), $|G|$ is,
$\left|G(\Delta)\right|^{2}=G_{1}^{2}G_{2}^{2}\left[x^{2}+2x\cos(\Delta-\phi_{T})+1\right].$ (9)
The signal-to-noise ratio (SNR) has a crucial role in sensing include imaging. In this section we calculate noise term of the system which arises from quantum fluctuations. To this end we need to calculate $\langle\Delta S^{2}\rangle_{\Psi}=\langle\Psi|S^{2}|\Psi\rangle$. Hence the uncertainty in signal measurement is given by,
$\langle\Delta S^{2}\rangle=\beta^{2}\left(2|G|^{2}-1\right).$ (10)
Interestingly, the quantum uncertainty given by Eq. (10) is related to the optical properties of the object through G parameter. Therefore, the image can be also reconstructed using quantum noise (see section V.1). The SNR (defined by $\langle S\rangle^{2}/\langle\Delta S^{2}\rangle$) can be calculated as:
$\text{SNR}=4\alpha^{2}\Gamma\cos^{2}(\phi_{G}+\delta).$ (11)
Here the factor $\Gamma$ is defined as $\Gamma=|G|^{2}\left(2|G|^{2}-1\right)^{-1}$. Since $|G|^{2}=1+|g|^{2}+|r|^{2}\geq 1$ is satisfied (see Sec. II), $\Gamma$ has a finite value.
@@ IV Description Characteristic of the System
In the following for better illustration, the spacial limits $x\rightarrow 0$ (opaque object) and $x\rightarrow 1$ (transparent object) are demonstrated, separately.
IV.1 The Limit $x\rightarrow 0$
For an object with low transmission co-efficient the condition $x\rightarrow 0$ is satisfied. In this case, $|G|$ tends to $G_{1}G_{2}$ and $\phi_{G}\approx x\sin(\Delta-\phi_{T})$. Substituting these results in Eq. (7) and ignoring $O(x^{3})$ and higher terms, one finds,
@@ IV.1 The Limit $x\rightarrow 0$
The small change in $\phi_{T}$ leads to the variation in the output signal as follows,
@@
For the leading purposes, we investigate the last equation for $\delta=k\pi/2$, where $k$ is an integer number.
@@
This equation reveals that, for an object with low transmission, the small change in the phase leads to a change in output signal which is proportional to $G_{1}G_{2}$. Therefore sensitivity of a fixed system can be enhanced by nonlinear process gains which is depend on tunable parameters such as the phase and amplitude of the pump field. From this equation, it is also concluded that behavior of the system is depend on selection of $k$ number such that for the odd values of $k$, output signal related on $x$ while for an even value, this quantity is proportional to $x^{2}$. In the limit of $x\rightarrow 0$, Eq. (11) becomes,
$\text{SNR}=\begin{cases}4\alpha^{2}\Gamma\phi_{G}^{2}&k\,\,\text{is odd}\\ 4\alpha^{2}\Gamma&k\,\,\text{is even}\end{cases},$ (15)
where, $\lim_{{}_{x}\rightarrow 0}\Gamma=G_{1}G_{2}\left(2G_{1}G_{2}-1\right)^{-1}$. In the low and high gain regime $\Gamma$ approaches to $1$ and $0.5$, respectively. Finally, from Eq. (14), one can conclude that SNR depend on $x^{2}$ for odd values of $k$ while for even values SNR is independent of $x$. Eq. (15) implies that SNR can be improved by adjusting the mean photon number in the input signal field of NOPO.
IV.2 The Limit $x\rightarrow 1$
The limit $x\rightarrow 1$ is achieved in high gain regime and high transmission. In this regime $\phi_{G}=(\Delta-\phi_{T})/2$ and Eq.(9) takes form,

### v2

@@ Abstract
In this work a quantum imaging setup based on undetected squeezed photons is employed for metrological applications such as sensitive phase measurement and quantum imaging. Unlike traditional quantum imaging with undetected photons, as introduced by A. Zeilinger et al., the proposed setup incorporates a homodyne detection system and enhances the brightness of the quantum light using optical parametric oscillators (OPOs). The inclusion of OPOs may challenge the validity of the low-gain approximation, necessitating the development of a new theoretical framework that extends beyond this approximation. The results demonstrate a higher signal-to-noise ratio, which serves as a key metric for both image quality and phase-measurement accuracy. Furthermore, an imaging protocol is introduced to effectively suppress background noise. Notably, this protocol enables the extraction of image information encoded in quantum fluctuations (noise), paving the way for non-disruptive imaging. This is particularly significant in the field of bio-imaging, where imaging sensitive living cells with a low damage threshold is of critical importance.
Interferometry provides Interferometry serves as a powerful technique to measure for measuring physical quantities with exceptional precision and sensitivity[1]. The interferometry is used in direct detection of gravitational waves [2]. The sensitivity of the conventional interferometers is bounded by Shot Noise Limit (SNL) [3, 4, 5]. One of the primary goals in interferometry is to overcome SNL. Using non-classical properties of lights provides promising resources to achieve this [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Recent advancements in nonlinear optics have spurred interest in nonlinear interferometry to achieve highly sensitive and accurate interferometers. [17, 18, 19, 20]. In particular, SU(1,1) interferometers, which are constructed by replacing the beam splitters with two mode squeezer devices (e.g Optical Parametric Amplifiers (OPAs)), offer sensitivity enhancement beyond SNL [21, 17, 22, 23, 24, 25, 26, 27, 28]. Non-degenerate SU(1,1) interferometers operate at distinct wavelengths for interaction and detection parties (e.g. mid-infrared for interaction party and optical spectrum for detection) [29, 30, 31, 32]. More precisely, these devices generate two-mode squeezed light at different wavelengths (sinal and idler beams) which enables quantum metrological applications in spectroscopy [33, 34], imaging [35, 36], optical coherent tomography(OCT)[3, 37] and bio-imaging of living cells [38, 39].
Quantum Imaging with Undetected Photon (QUIP), introduced by Zeilinger et al., utilizes an SU(1,1) interferometer. In this scheme, the idler beam interacts with an object, while the image is formed by detecting the signal beams, which never interacts with the object[40, 31]. In the original setup, a signal-idler photon pair is generated via spontaneous parametric down-conversion (SPDC) in a nonlinear crystal. The idler beam interacts with the object and is subsequently combined with a second idler beam produced in an identical nonlinear crystal. This combination induces interference between the two corresponding signal beams emitted from the respective crystals. Crucially, the observed interference patterns in the signal beams stem from the path indistinguishably (or identity) of the idler beams, rather than any mechanism related to induced emission [41]. Theoretical analyses of QIUP typically assume operation in the low-gain regime [42]. However, this approximation restricts the framework applicability to high-gain systems. In this work, we develop a generalized theoretical model for QIUP that transcends the low-gain assumption. To achieve this, we solve the system dynamics in the Heisenberg picture (as opposed to the Schrodinger picture) and derive analytical expressions for the output creation and annihilation operators at the device’s output port, explicitly incorporating the classical pump field. To improve imaging quality, quantified by the Signal-to-Noise Ratio (SNR), and to investigate the performance of QIUP in high-gain regimes, each nonlinear crystal is placed within a con-focal cavity. This configuration replaces the crystals with two identical Non-degenerate Optical Parametric Oscillators (NOPOs). In each NOPO generate a signal-idler biphoton state via cavity-PDC process[43], which exhibits significant squeezing, enabling SNR enhancement by leveraging squeezing as a quantum resource. Consequently, we utilize homodyne detection—rather than EMCCD—for data acquisition. We propose an imaging protocol designed to suppress noise. Additionally, we introduce a novel method for constructing images from objects with low damage thresholds, termed Quadrature-Noise Shadow Imaging (QSI). This scheme relies on quadrature quantum noise measurements [44]. QSI permits imaging with low photon flux while eliminating the detrimental dark-count noise of sensors in practical setups [44, 45, 46]. Intriguingly, we demonstrate that image information is encoded into the noise itself, enabling reconstruction via QSI.
@@
This paper is organized as follow: In the next section the setup under consideration is described. Section 2 is devoted to the theoretical description of the model. In first subsection of Sec.3 opaque objects are characterized while in second subsection transparent objects is addressed. Accordingly two imaging protocols are introduced in Sec. 4. Finally the conclusion and remarks of this work are provided in Sec. 5.
1 The Setup
@@
This section demonstrates the proposed setup in detail. The schematic diagram of the setup is depicted in figure 1. The imaging process in this figure can be explained through four stages. In the first stage the two mode squeezed state beams, including idler (red) and signal(green) beams, are generated in the output of the first NOPO. Then a dichroic mirror $DM1$ separates them into different paths: The signal (green) beam arrives to the second NOPO and the idler (red) beam is reflected through $DM1$ toward the object. The idler interacts with the object which is modeled by a lossless beam splitter with transmission $Te^{i\phi_{T}}$ (reflection $Re^{i\phi_{R}}$). In order to provide path identity, a portion of the idler, which is transmitted through the object, is aligned with second idler beam, generated in the second NOPO. Finally, the information about the object is extracted from signal beam due to the quantum interference of idler states. Since the object information are encoded in the signal quadratures, a balanced homodyne detection is applied to the outcome signal of the second NOPO to extract the image information. It is worth noting that undetected light refers to the idler beam as no measurement applied to it.
. Refer to caption
@@
In the next section, we present a theoretical framework describing how image information is extracted.
2 Theoretical Description of The Model
Considering a local oscillator prepared in coherent state, $|\beta e^{i\phi_{\beta}}\rangle$, the output electrical signal of the balanced homodyne detection (depicted in Fig. 1) can be calculated as:
@@ 2 Theoretical Description of The Model
The annihilation (creation) operator $a_{S2}$ ($a_{S2}^{\dagger}$) is correspond to the output signal beam from second NOPO. The image information is encoded in this beam, Assuming that the NOPOs are pumped by classical fields with fixed phase, $\phi_{P1}$ and $\phi_{P2}$, respectively, we can write:
@@
$a_{IT}=Te^{i\phi_{T}}a_{I1}+Re^{i\phi_{R}}a_{IL},$ (3)
here, $a_{IL}$ is relate to the empty port of the beam splitter (object). It is straightforward to show that
@@
Two beams ($a_{S1}$ and $a_{IT}$) are then mixed in the second NOPO to built up beam $S_{2}$:
@@
where $G_{2}$, $g_{2}$ and $\phi_{P2}$ are the gain parameters and pump phase of the second crystal. Substituting Eqs. (2a) and (4) into Eq. (5) leads to,
$a_{S2}=Ga_{S0}+ga_{I0}^{\dagger}+ra_{IL}^{\dagger},$ (6)
@@
with $G=G_{1}G_{2}+g_{1}g_{2}T\exp\left(i\phi_{1}\right)$, $g=g_{1}G_{2}\exp(i\phi_{P1})+G_{1}g_{2}T\exp\left(i\phi_{2}\right)$ and $r=g_{2}R\exp(i\phi_{3})$. Here, the parameters $\phi_{1}=\phi_{P2}-\phi_{P1}-\phi_{T}$, $\phi_{2}=\phi_{P2}-\phi_{T}$ and $\phi_{3}=\phi_{P2}-\phi_{R}$ are defined for simplicity. It is worth noting, since $|G|^{2}-|g|^{2}-|r|^{2}=1$ is satisfied, the commutation relation $\left[a_{S2},a_{S2}^{\dagger}\right]=1$ is also preserved.
3 System Characteristics
If the input signal beam is in coherent state $|\alpha e^{i\phi_{\alpha}}\rangle$ and all other modes are in the vacuum state then the expectation value of the output signals becomes:
$\langle S(\delta,\Delta)\rangle=\,\,\,_{IL}\langle 0|_{I0}\langle 0|_{S0}\langle\alpha e^{i\phi_{\alpha}}|\,\,S(\delta,\Delta)\,\,|\alpha e^{i\phi_{\alpha}}\rangle_{S0}|0\rangle_{I0}|0\rangle_{IL}=\frac{1}{2}\alpha\beta|G|\cos(\phi_{G}(\Delta)+\delta),$ (7)
where, $\delta=\phi_{\alpha}+\phi_{\beta}$, $\Delta=\phi_{P2}-\phi_{P1}$, $\phi_{G}$ and magnitude $|G|$ is defined as,
$\phi_{G}(\Delta)=\sin^{-1}(\frac{Tg_{1}g_{2}\sin(\Delta-\phi_{T})}{|G|}),$ (8)
and
$|G(\Delta)|^{2}=G_{1}^{2}G_{2}^{2}\left[x^{2}+2x\cos(\Delta-\phi_{T})+1\right],$ (9)
with $x=Tg_{1}g_{2}(G_{1}G_{2})^{-1}$. Figures 2-a and 2-b compare the visibility of homodyne detection and conventional Photon Number Counting(PNC).
Figure 2: Visibility of the image, $\frac{g_{1}g_{2}}{G_{1}G_{2}}=0.25$ (dot curve), $\frac{g_{1}g_{2}}{G_{1}G_{2}}=0.5$ (dot-dashed curve), $\frac{g_{1}g_{2}}{G_{1}G_{2}}=0.75$ (dashed curve) and $\frac{g_{1}g_{2}}{G_{1}G_{2}}=1$ (solid curve) (a) homodyne detection which is defined by $V=\frac{\max(S^{rms})-\min(S^{rms})}{\max(S^{rms})+\min(S^{rms})}$, where $S^{rms}$ denotes root mean square of output signal. (b) Conventional photon counting which is defined by $V=\frac{\max(n_{S2}^{rms})-\min(n_{S2}^{rms})}{\max(n_{S2}^{rms})+\min(n_{S2}^{rms})}$ The root mean square of the photon number is presented by $n_{S2}^{rms}$.
These figures reveal that the visibility of homodyne detection is a linear function of the object’s transmitivity, $T$ and PNC visibility is linear only for low gain regime. The nonlinear response of PNC for high gain regime implies distortions that degrade the quality of the recorded image, This fact is advantageous for roubust performance for homodyne detection across a range of, $T$.
The signal-to-noise ratio (SNR) is a crucial role in metrology include imaging. SNR is defined as follow:
$SNR=\frac{\langle S^{2}\rangle}{\langle\Delta S^{2}\rangle},$ (10)
here $\langle S^{2}\rangle=\frac{\beta^{2}}{16}\left[4|G|^{2}\alpha^{2}\cos^{2}\left(\phi_{G}+\delta\right)+2|G|^{2}-1\right]$ is the strength of the signal and $\langle\Delta S^{2}\rangle=\frac{\beta^{2}}{16}\left(2|G|^{2}-1\right)$ denotes its variance raised by quantum fluctuations. Interestingly, $\langle\Delta S^{2}\rangle$ is related to the optical properties of the object via parameter, G. This fact brings up idea of image construction from noise in our mind. For the case of $\alpha\gg 1$, both homodyne and PNC detection scheme expression of SNR can be simplified to:
$\text{SNR}_{HD}=\frac{4\alpha^{2}|G|^{2}}{\left(2|G|^{2}-1\right)}\cos^{2}(\phi_{G}+\delta),$ (11)
$\text{SNR}_{PNC}=\frac{\alpha^{2}|G|^{2}}{\left(2|G|^{2}-1\right)}.$ (12)
$\text{SNR}_{HD}$ is depicted in 3 as a function of $\Delta-\phi_{T}$. Figure 3 depicts $\text{SNR}_{HD}$ as a function of $\Delta-\phi_{T}$ for the fixed values of the gain. The $\text{SNR}_{HD}$ curve is symmetric around its peak value which occurs at $\Delta-\phi_{T}=\pi$. Variation of the $\text{SNR}_{HD}$ becomes smoother as the object becomes opaquer.
The ratio $\text{SNR}_{HD}/\text{SNR}_{PC}$ is scaled by a factor of $4\cos^{2}(\phi_{G}+\delta)$. This quantity is plotted in Fig. 4. This figure reveals that $\cos^{2}(\phi_{G}+\delta)$ remains constant for the case of $x\ll 1$. This finding indicates that, $\text{SNR}_{HD}$ can be improved by a factor of order 4 with respect to $SNR_{PNC}$ for transparent objects.
Figure 3: homodyne detection SNR as a function of $\Delta-\phi_{T}$ with $\alpha=100$. (a) $\frac{g_{1}g_{2}}{G_{1}G_{2}}=0.02$, (b) $\frac{g_{1}g_{2}}{G_{1}G_{2}}=0.17$, (c) $\frac{g_{1}g_{2}}{G_{1}G_{2}}=0.75$, (d) $\frac{g_{1}g_{2}}{G_{1}G_{2}}=0.96$
Figure 4: Contour plot of $\cos^{2}(\phi_{G}+\delta)$ (a)$\delta=0$ and (b) $\delta=\pi/2$.
@@ 3 System Characteristics
In the following for better illustration, we discuss more about two extremes of $x\rightarrow 0$ (opaque object) and $x\rightarrow 1$ (transparent object), separately.
3.1 Opaque Object, $x\rightarrow 0$
For transparent object, where $x\rightarrow 0$, $|G|\rightarrow G_{1}G_{2}$ and $\phi_{G}\approx x\sin(\Delta-\phi_{T})$. Retaining tterms up to second order in $x$, the expectation value of $S$ becomes:
@@ 3.1 Opaque Object, $x\rightarrow 0$
A small change in $\phi_{T}$ leads to the variation:
@@
For $\delta=k\pi/2$, (with integer $k$),.
@@
This equation reveals that, in low transmission regime, the small change in the phase leads to a change in output signal which is proportional to $G_{1}G_{2}$. Therefore sensitivity of a fixed system can be enhanced by nonlinear process gains which is depend on involved parameters such as the phase and amplitude of the pump field. It is also concluded that behavior of the system in this regime is depend on the parity of $k$, such that for the odd values of $k$, output signal scaled by $x$ while for an even value, this quantity is proportional to $x^{2}$. Hence, in this limit we can write:
$\text{SNR}_{HD}=\begin{cases}4\alpha^{2}\Gamma\phi_{G}^{2}&k\,\,\text{is odd}\\ 4\alpha^{2}\Gamma&k\,\,\text{is even}\end{cases},$ (16)
where, $\Gamma=\lim_{x\rightarrow 0}(G_{1}G_{2}(2G_{1}G_{2}-1)^{-1})$ , approaching 1 in the low-gain regime and 0.5 in the high-gain regime. $\text{SNR}_{HD}$ scales as $x^{2}$ for odd $k$ and for even $k$ it is independent of $x$. Eq. (16) implies that $\text{SNR}_{HD}$ can be improved by adjusting the mean photon number in the input signal field of NOPO.
3.2 High gain, High Transmission, $x\rightarrow 1$
In the high gain, high transmission regime ($x\rightarrow 1$),we obtain $\phi_{G}=(\Delta-\phi_{T})/2$ and: